ganeshie8 (ganeshie8):

\(\gcd((p+1)/2,~p-1) = ?\) \(p\) is an odd prime

OpenStudy (dan815):

what is the GCD algorithm

ganeshie8 (ganeshie8):

wolfram says the answer is always either 2 or 1 http://www.wolframalpha.com/input/?i=Table%5Bgcd%28%28p%2B1%29%2F2%2C+p-1%29%2C+%7Bp%2C3%2C13%7D%5D

ganeshie8 (ganeshie8):

gcd(a, b) = gcd(a-b, b) = gcd(a, b-a)

ganeshie8 (ganeshie8):

i have been messing with that but no success..

OpenStudy (dan815):

1/2, 2, 1/2, 1

OpenStudy (dan815):

okay (p+1)/2 = p/2 + 1/2, and p-1 = 2*(p/2 - 1/2)

ganeshie8 (ganeshie8):



OpenStudy (dan815):

i think showing this pattern recurring is better

OpenStudy (dan815):

we can show that all oodds will be 1 and 2 repeating

OpenStudy (dan815):

let p=2n+1

OpenStudy (dan815):

GCD((2n+1)+1)/2,2n-1)

ganeshie8 (ganeshie8):

\(p\) is given as an odd prime so..

OpenStudy (dan815):

if we show this happens to all odd numbers then we can atleast say it happens to all odd primes, the patternn of the 1s and 2s might be random

OpenStudy (dan815):

GCD((2n+1)+1)/2,2n-1) = 1 or 2 in fact we can evn do 1 step better by changing n=2k and 2k+1

OpenStudy (dan815):

to show every 2nd add is 2 and every 1st odd is 1

ganeshie8 (ganeshie8):

Oh! for odd numbers also it is giving 1 or 2 didn't notice before.. interesting..

OpenStudy (dan815):

right

OpenStudy (dan815):

oops i mean (2n-1)-1

OpenStudy (dan815):